# On the Necessity of Collaboration in Online Model Selection with Decentralized Data

### v2

@@ 4.2.2 Updating hypotheses
{¯$\left\{\begin{split}\bar{\nabla}_{t,i}=&\frac{1}{M}\sum^{M}_{j=1}\tilde{\nabla}^{(j)}_{t,i},\\ \nabla_{\bar{{\bf w}}_{t+1,i}}\psi_{t,i}(\bar{{\bf w}}_{t+1,i})=&\nabla_{{\bf w}_{t,i}}\psi_{t,i}({\bf w}_{t,i})-\bar{\nabla}_{t,i}\\ {\bf w}_{t+1,i}=&\mathop{\arg\min}_{{\bf w}\in\mathcal{F}_{i}}\mathcal{D}_{\psi_{t,i}}({\bf w},\bar{{\bf w}}_{t+1,i}),\\ \psi_{t,i}({\bf w})=&\frac{1}{2\lambda_{t,i}}\cdot\|{\bf w}\|^{2}_{\mathcal{F}_{i}},\end{split}\right.$ (12)
@@ 4.3 Regret bounds
To obtain high-probability regret bounds that adapt to the complexity of individual hypothesis space, we first give a new Bernstein’s inequality for martingale.
@@
Let $X_{1},\ldots,X_{n}$ be a bounded martingale difference sequence w.r.t. the filtration $\mathcal{H}=(\mathcal{H}_{k})_{1\leq k\leq n}$ and with $|X_{k}|\leq a$. Let $Z_{t}=\sum^{t}_{k=1}X_{k}$ be the associated martingale. Denote the sum of the conditional variances by $\Sigma^{2}_{n}=\sum^{n}_{k=1}\mathbb{E}\left[X^{2}_{k}|\mathcal{H}_{k-1}\right]\leq v,$ where $v\in[0,B]$ is a random variable and $B\geq 2$ is a constant. Then for all constants $a>0$, with probability at least $1-2\lceil\log{B}\rceil\delta$,
@@
$\lambda_{t,i}=$ $\frac{U_{i}}{2G_{i}\sqrt{\left(1+\frac{K-J}{(J-1)M}\right)\cdot\max\left\{\frac{(K-J)^{2}}{(J-1)^{2}},t\right\}}}.$
@@ 4.5 Answers to Question 1
Assuming that $5\leq K\leq\min\{d,T\}$. For each $i\in[K]$, let $\mathcal{F}_{i}=\{f_{i}({\bf x})=r_{i}{\bf e}^{\top}_{i}{\bf x}\}$ and $\mathcal{D}_{i}=[\min_{{\bf x}\in\mathcal{X}}f_{i}({\bf x}),\max_{{\bf x}\in\mathcal{X}}f_{i}({\bf x})]$, where ${\bf e}_{i}$ is the standard basis vector in $\mathbb{R}^{d}$ and $0\leq r_{i}\leq 1$. Denote by $\sup$ the supremum over all examples.
(i) There are no additional constraints on the problem. Let $\ell(v,y)=|v-y|$. The regret of any algorithm satisfies: $\lim_{T\rightarrow\infty}\sup\max_{i\in[K]}\mathrm{Reg}_{D}(\mathcal{F}_{i})\geq 0.25M\sqrt{T\ln{K}}$;
(ii) The per-round time complexity on each client is limited to $O(J)$. Let $\ell(v,y)=1-v\cdot y$. The regret of any noncooperative algorithm with outputs in $\cup_{i\in[K]}\mathcal{D}_{i}$ satisfies: with probability at least $1-\delta$, $\sup\max_{i\in[K]}\mathrm{Reg}_{D}(\mathcal{F}_{i})\geq cM\sqrt{KTJ^{-1}}+M\sqrt{0.5T\ln{\delta^{-1}}}$, where $c>0$ is a constant.
The assumption that the outputs of any noncooperative algorithm belong to $\cup_{i=\in[K]}\mathcal{D}_{i}$ is natural, and can be removed in the case of $J=1$.
Next we define a noncooperative algorithm named NCO-OMS.
@@
$\nabla_{\bar{{\bf w}}_{t+1,i}}\psi_{t,i}(\bar{{\bf w}}_{t+1,i})=$ $\nabla_{{\bf w}^{(j)}_{t,i}}\psi_{t,i}\left({\bf w}^{(j)}_{t,i}\right)-\tilde{\nabla}^{(j)}_{t,i},\quad{\bf w}^{(j)}_{t+1,i}=\mathop{\arg\min}_{{\bf w}\in\mathcal{F}_{i}}\mathcal{D}_{\psi_{t,i}}({\bf w},\bar{{\bf w}}_{t+1,i}).$
@@
The definitions of $\tilde{{\bf c}}^{(j)}_{t}$ and ~$\tilde{\nabla}^{(j)}_{t,i}$ follow FOMD-OMS. It is worth mentioning that any non-cooperative algorithm can be constructed by independently running any existing algorithm that solves centralized online model selection and satisfies the computational constraints. It is easy to show the regret of NCO-OMS satisfies
$\forall i\in[K],~{}\mathrm{Reg}_{D}$ $(\mathcal{F}_{i})=O\left(M\left(B_{i,1}\sqrt{\left(1+g_{K,J}\right)T}+B_{i,2}g_{K,J}\ln\frac{1}{\delta}+B_{i,3}\sqrt{g_{K,J}T\ln\frac{1}{\delta}}\right)\right).$
@@
Case 1: $J=\Theta(K)$. Collaboration is unnecessary.
In this case, there are no additional constraints on clients. Both FOMD-OMS and NCO-OMS enjoy a $O(MB_{i,1}\sqrt{T\log(KT)})$ regret (see Theorem 3). By Theorem 4, FOMD-OMS and NCO-OMS are optimal in terms of the dependence on $M$ and $T$. Thus collaboration is unnecessary.
Case 2: $J=o(K)$. Collaboration is necessary.
In this case, the time complexity on clients is limited to $o(K)$. FOMD-OMS enjoys a $\tilde{O}(MB_{i,1}\sqrt{T}+B_{i,3}\sqrt{MKTJ^{-1}})$ regret. By the second lower bound in Theorem 4, the upper bound of FOMD-OMS is smaller than the lower bound of any noncooperative algorithm. Thus collaboration is necessary.
@@ 6 Application
We defer the algorithm in the appendix. Let $R=T$ and $J=2$. Then we obtain Table 2. According to Section 4.4, FOMD-OMS enjoys a $O(D)$ per-round time complexity on each client.
Next we compare FOMD-OMS with vM-KOFL, eM-KOFL [22] and POF-MKL [23]. Table 3 gives the regret bounds and download cost of the three algorithms. The per-round time complexity of the three algorithms is $O(KD)$. Recalling the answer to Question 1 (see Section 4.5), collaboration is unnecessary.
FOMD-OMS is better than the three algorithms. (i) The regret bounds of the three algorithms can not adapt to the complexity of the optimal hypothesis space $\mathcal{F}_{i^{\ast}}$. (ii) FOMD-OMS has a better dependence on $M$ than POF-MKL. (iii) In the case of $K\leq M$, FOMD-OMS enjoys a similar regret bound with vM-KOFL and eM-KOFL at a smaller download cost or computational cost.
Table 3: Regret bound and download cost.
Algorithm Regret bound download
vM-KOFL $\tilde{O}\left(\mathfrak{C}M\sqrt{T\ln{K}}+\mathfrak{C}_{i}\frac{MT}{\sqrt{D}}\right)$ $O(DKM)$
eM-KOFL $\tilde{O}\left(\mathfrak{C}M\sqrt{T\ln{K}}+\mathfrak{C}_{i}\frac{MT}{\sqrt{D}}\right)$ $O(DM)$
POF-MKL $\tilde{O}\left(\mathfrak{C}M\sqrt{KT}+\mathfrak{C}_{i}\frac{MT}{\sqrt{D}}\right)$ $O(DKM)$
@@ 7.1 Experimental setting
$\forall t\in[T],~{}\eta_{t}=\max\left\{\frac{\sqrt{\ln{(KT)}}}{2\sqrt{\left(K-1\right)T}},\frac{1}{2(K-2)}\right\},~{}\lambda_{t,i}=\frac{U_{i}}{2G_{i}\sqrt{\left(K-1\right)\cdot\max\left\{(K-2)^{2},t\right\}}}.$
@@
Table 4: Basic information of datasets. #num is the number of examples. #fea is the number of features.
  Dataset #num #fea Dataset #num #fea Dataset # num #fea Dataset #num #fea
elevators 16590 18 bank 8190 32 TomsHardware 28170 96 Twitter 50000 77
ailerons 13750 40 calhousing 14000 8 Year 51630 90 Slice 53500 384
@@ 8 Conclusion
In this paper, we have studied the necessity of collaboration in OMS-DecD from the perspective of computational constraints. We demonstrate that collaboration is unnecessary when there are no additional constrains on the problem, while it becomes necessary if the time complexity on each client is limited to $o(K)$. Our work clarifies the unnecessary nature of collaboration in previous algorithms for the first time, gives conditions under which collaboration is necessary, and provides inspirations for studying the problem from constraints beyond computational constrains.

### v4

@@ 5.2.2 Updating hypotheses
{¯$\left\{\begin{split}\bar{\nabla}_{t,i}=&\frac{1}{M}\sum^{M}_{j=1}\tilde{\nabla}^{(j)}_{t,i},\\ \nabla_{\bar{{\bf w}}_{t+1,i}}\psi_{t,i}(\bar{{\bf w}}_{t+1,i})=&\nabla_{{\bf w}_{t,i}}\psi_{t,i}({\bf w}_{t,i})-\bar{\nabla}_{t,i},\\ {\bf w}_{t+1,i}=&\mathop{\arg\min}_{{\bf w}\in\mathcal{F}_{i}}\mathcal{D}_{\psi_{t,i}}({\bf w},\bar{{\bf w}}_{t+1,i}),\\ \psi_{t,i}({\bf w})=&\frac{1}{2\lambda_{t,i}}\cdot\|{\bf w}\|^{2}_{\mathcal{F}_{i}},\end{split}\right.$ (12)
@@ 5.3 Regret bounds
To obtain high-probability regret bounds that adapt to the complexity of individual hypothesis space, we establish a new Bernstein’s inequality for martingale.
@@
Let $X_{1},\ldots,X_{n}$ be a bounded martingale difference sequence w.r.t. the filtration $\mathcal{H}=(\mathcal{H}_{k})_{1\leq k\leq n}$ and with $|X_{k}|\leq a$. Let $Z_{t}=\sum^{t}_{k=1}X_{k}$ be the associated martingale. Denote the sum of the conditional variances by $\Sigma^{2}_{n}=\sum^{n}_{k=1}\mathbb{E}\left[X^{2}_{k}|\mathcal{H}_{k-1}\right]\leq v,$ where $v\in[0,B]$ is a random variable and $B\geq 2$ is a constant. Then for any constant $a>0$, with probability at least $1-2\lceil\log{B}\rceil\delta$,
@@
$\lambda_{t,i}=$ $\frac{U_{i}}{2G_{i}\sqrt{\left(1+\frac{K-J}{(J-1)M}\right)\cdot\left(\frac{(K-J)^{2}}{(J-1)^{2}}\vee t\right)}}.$
@@ 5.5 Answers to Question 1
Assuming that $5\leq K\leq\min\{d,T\}$. For each $i\in[K]$, let $\mathcal{F}_{i}=\{f_{i}({\bf x})={\bf e}^{\top}_{i}{\bf x}\}$ and $\mathcal{D}_{i}=[\min_{{\bf x}\in\mathcal{X}}f_{i}({\bf x}),\max_{{\bf x}\in\mathcal{X}}f_{i}({\bf x})]$, where ${\bf e}_{i}$ is the standard basis vector in $\mathbb{R}^{d}$. Denote by $\sup$ the supremum over all examples.
(i) There are no computational constraints on clients. Let $\ell(v,y)=|v-y|$. The regret of any algorithm for OMS-DecD satisfies: $\lim_{T\rightarrow\infty}\sup\max_{i\in[K]}\mathrm{Reg}_{D}(\mathcal{F}_{i})\geq 0.25M\sqrt{T\ln{K}}$;
(ii) The per-round time complexity on each client is limited to $O(J)$. Let $\ell(v,y)=1-v\cdot y$. The regret of any, possibly randomized, noncooperative algorithm with outputs in $\cup_{i\in[K]}\mathcal{D}_{i}$ satisfies: with probability at least $1-\delta$, $\sup\mathbb{E}[\max_{i\in[K]}\mathrm{Reg}_{D}(\mathcal{F}_{i})]\geq 0.1M\sqrt{KTJ^{-1}}+M\sqrt{0.5T\ln{(M/\delta)}}$, where the expectation is taken over the randomization of algorithm.
The assumption that the outputs of any noncooperative algorithm belong to $\cup_{i=\in[K]}\mathcal{D}_{i}$ is natural, and can be removed in the case of $J=1$. Next we define a noncooperative algorithm, NCO-OMS.
@@
$\nabla_{\bar{{\bf w}}_{t+1,i}}\psi_{t,i}(\bar{{\bf w}}_{t+1,i})=$ $\nabla_{{\bf w}^{(j)}_{t,i}}\psi_{t,i}\left({\bf w}^{(j)}_{t,i}\right)-\tilde{\nabla}^{(j)}_{t,i},\quad{\bf w}^{(j)}_{t+1,i}=\mathop{\arg\min}_{{\bf w}\in\mathcal{F}_{i}}\mathcal{D}_{\psi_{t,i}}({\bf w},\bar{{\bf w}}_{t+1,i}),$
where the definitions of $\tilde{{\bf c}}^{(j)}_{t}$ and ~$\tilde{\nabla}^{(j)}_{t,i}$ follow FOMD-OMS.
@@
The pseudo-code of NCO-OMS is shown in Algorithm 4. It is easy to prove the regret of NCO-OMS satisfies: with probability at least $1-\Theta\left(M\log(KT)\right)\cdot\delta$,
$\forall i\in[K],~{}\mathrm{Reg}_{D}$ $(\mathcal{F}_{i})=O\left(M\left(B_{i,1}\sqrt{\left(1+g_{K,J}\right)T}+B_{i,2}g_{K,J}\ln\frac{1}{\delta}+B_{i,3}\sqrt{g_{K,J}T\ln\frac{1}{\delta}}\right)\right),$
@@
Case 1: There are no computational constraints on clients. Collaboration is unnecessary.
Let $J=\Theta(K)$ in FOMD-OMS and NCO-OMS. By Theorem 3, both FOMD-OMS and NCO-OMS enjoy a $O(MU_{i}G_{i}\sqrt{T}+MC_{i}\sqrt{T\ln(KT)})$ regret. By Theorem 4, FOMD-OMS and NCO-OMS are nearly optimal in terms of the dependence on $M$ and $T$. Thus collaboration is unnecessary.
Case 2: The per-round time complexity on each client is limited to $o(K)$. Collaboration is necessary.
Let $J=o(K)$ in FOMD-OMS and Theorem 4. By Theorem 3, FOMD-OMS enjoys a $O(MB_{i,1}\sqrt{T}+B_{i,3}\sqrt{MKTJ^{-1}\ln{\delta^{-1}}})$ regret, which is smaller than the lower bound on the regret of any noncooperative algorithm (see Theorem 4). Thus collaboration is necessary.
@@ 7 Application to Distributed OMKL
We defer the algorithm in appendix. Let $R=T$ and $J=2$. We compare FOMD-OMS with eM-KOFL [23] and POF-MKL [24]. Table 3 gives the results.
We observe that FOMD-OMS significantly improves the computational complexity of eM-KOFL and POF-MKL (by a factor of $O(K)$). The per-round time complexity of the two algorithms is $O(DK)$. Recalling the answer to Question 1 (see Section 5.5), collaboration in eM-KOFL and POF-MKL is unnecessary.
Next we compare the regret bound of the three algorithms. Recalling that $\mathfrak{C}_{i}\leq\mathfrak{C}$. The regret bounds of eM-KOFL and POF-MKL can not adapt to the complexity of individual hypothesis space. (i) The regret bound of FOMD-OMS is better than that of POF-MKL in relation to its dependence on $M$ and $\mathfrak{C}_{i}$. (ii) In the case of $K=O\left(\frac{\mathfrak{C}}{\mathfrak{C}_{i}}M\cdot\ln{K}\right)$, the regret bound of FOMD-OMS is better than that of eM-KOFL. (iii) In the case of $K=\Omega\left(\frac{\mathfrak{C}}{\mathfrak{C}_{i}}M\cdot\ln{K}\right)$, the regret bound of FOMD-OMS is worse than that of eM-KOFL. If $K$ is sufficiently large, the regret bound of eM-KOFL is better than that of FOMD-OMS by a factor of $O(\sqrt{K})$.
Table 3: Comparison with previous algorithms. $D$ is the number of random features [26]. $\tilde{O}(\cdot)$ hides $O(\mathrm{poly}(\ln\frac{T}{\delta}))$. Time (s) is the per-round time complexity on client.
Algorithm Regret bound Time (s) Download (bits)
eM-KOFL $\tilde{O}\left(\mathfrak{C}M\sqrt{T\ln{K}}+\frac{\mathfrak{C}_{i}MT}{\sqrt{D}}\right)$ $O(DK)$ $O(DM\log{K})$
POF-MKL $\tilde{O}\left(\mathfrak{C}M\sqrt{KT}+\frac{\mathfrak{C}_{i}MT}{\sqrt{D}}\right)$ $O(DK)$ $O(DMK)$
FOMD-OMS $\tilde{O}\left(\mathfrak{C}_{i}M\sqrt{T\ln{K}}+\sqrt{\mathfrak{C}\mathfrak{C}_{i}MKT}+\frac{\mathfrak{C}_{i}MT}{\sqrt{D}}\right)$ $O(D)$ $O(DM\log{K})$
@@ 8.1 Experimental setting
$\forall t\in[T],\quad\eta_{t}=\frac{\sqrt{\ln{(KT)}}}{2\sqrt{\left(K-1\right)T}}\wedge\frac{1}{2(K-2)},\quad\lambda_{t,i}=\frac{U_{i}}{2G_{i}\sqrt{\left(K-1\right)\cdot\left((K-2)^{2}\vee t\right)}}.$
@@
Table 4: Basic information of datasets.
  Dataset Number of Instances Number of features
bank 8,190 32
ailerons 13,750 40
calhousing 14,000 8
elevators 16,590 18
TomsHardware 28,170 96
Twitter 50,000 77
Year 51,630 90
Slice 53,500 384
@@ 9 Conclusion
In this paper, we have studied the necessity of collaboration in OMS-DecD from the perspective of computational constraints. We demonstrate that collaboration is unnecessary when there are no computational constrains on clients, while it becomes necessary if the time complexity on each client is limited to $o(K)$. Our work clarifies the unnecessary nature of collaboration in previous algorithms for the first time, gives conditions under which collaboration is necessary, and provides inspirations for studying the problem from constraints beyond computational constrains.